# Screening in Hubbard models with long-range interactions

### v1

@@ I.2 Phase diagrams
Figure 1: (a) Phase diagram of the one-dimensional $1/r$-Hubbard model with nearest-neighbor interactions $V_{\rm NN}$. (b) Phase diagram of the one-dimensional $1/r$-Hubbard model with $1/r$-long-range interactions. Dots: estimate for the critical interaction, $\overline{U}_{\rm c}$, with error bounds; continuous lines: polynomial fits, see Sect. V; dotted line: Hartree-Fock (HF) result for the transition between metal and CDW insulator.
For dominant Coulomb interactions, the separation line between the Mott-Hubbard insulator and the CDW insulator is given by $V_{\rm NN}=U/2$ to leading order, with a stabilization tendency toward the Mott-Hubbard insulator in the next order in $1/U$. The separation line in Fig. 1a (dashed line for $U\gtrsim 1.94$) is given by $V_{\rm NN}=0.51U$, in agreement with the strong-coupling prediction.
In the absence of inter-site interactions, the Mott-Hubbard transition occurs at $U_{\rm c}(V=0)=1$ [20, 6, 22]. The inclusion of a repulsive nearest-neighbor interaction stabilizes the metallic phase because the additional repulsive interaction softens the two-particle scattering potential in position space.
@@
On the other hand, in the atomic limit, the CDW and Mott-Hubbard insulators are separated by $V_{\rm c}^{\rm CDW}(U)=U/(2\ln(2))\approx 0.721U$ for the $1/r$-Hubbard model with $1/r$-long-range interactions, see appendix; in Fig. 1b we show the dashed line $V_{\rm LR}=0.767U$. Thus, the renormalization factor in the atomic limit is given by $R^{\rm al}_{\rm Mott}=(1/2)/[1/(2\ln 2)]=\ln 2\approx 0.693$. In Fig. 2, the Mott critical lines $U_{\rm c}^{\rm Mott}(V)$ of the $1/r$-Hubbard models with nearest-neighbor and $1/r$-long-range interactions are then mapped onto each other using $V_{\rm NN}=R_{\rm Mott}V_{\rm LR}$, $R_{\rm Mott}=0.69$.
@@ II.2.3 Density-density correlation function and Luttinger parameter
We extract the Luttinger exponent $K(U,V)$ from the structure factor,
@@
where we denote the structure factor in the thermodynamic limit by $\tilde{C}^{\rm NN}(q;U,V)$. Using Eq. (30), the Luttinger exponent can be calculated numerically with very good accuracy [29]. The Luttinger parameter can be used to locate the metal-insulator transition in one spatial dimension.
@@ III.1 Ground-state energy
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
Figure 4: Ground-state energy per lattice site at half band-filling in the thermodynamic limit, $e_{0}(n=1;U,V)$, for the $1/r$-Hubbard model with $1/r$-long-range interactions (crosses) from the extrapolation to the thermodynamic limit in Fig. 3, and for the extended $1/r$-Hubbard model (open triangles, see Ref. [23]). The continuous line is the exact result for $V=0$, $e_{0}(n=1;U,V=0)$, see Eq. (32).
@@ III.2 Two-particle gap
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
In the bare $1/r$-Hubbard model, the Mott transition is solely governed by the on-site interaction. Once the Hubbard interaction exceeds the bandwidth, $U>W$, electrons become ‘localized’, and the metallic state goes over to a Mott insulator. In the presence of inter-site interactions, the Mott transition is shifted to larger values which implies a more stable metallic Luttinger liquid. At first glance, the shift of the Mott transition to higher values in models with repulsive inter-site interactions seems counterintuitive. From a wave-mechanical viewpoint, however, the repulsive interaction between lattice sites softens the two-particle scattering potential. Figuratively speaking, particles scattered by the weaker interaction $V$ do not fully experience the stronger on-site interaction $U$. This ties in with the observation that the gap in the $1/r$-Hubbard model with $1/r$-long-range interactions is smaller than that for the extended $1/r$-Hubbard model for the same $v$ because the long-range parts of the inter-site interaction further smoothen the two-particle scattering potential in position space.
III.3 Momentum distribution
These considerations show that the Luttinger parameter must deviate from unity, $K(U,V)<1$, for all $(U,V)$, even though corrections to unity are (exponentially) small for $U,V\ll 1$. Therefore, the momentum distribution is a continuous function in the (extended) $1/r$-Hubbard model for all $U,V>0$.
III.4 Structure factor and CDW order parameter
@@
$\tilde{C}^{\rm NN}(q,n=1;U=0,V=0)=\frac{|q|}{\pi}\;,$ (40)
@@ IV.3 Luttinger parameter
(a) Refer to caption
(b) Refer to caption
As seen from Table 1, the critical value from the Luttinger parameter systematically overestimates the correct interaction strengths by some five percent. The same overshooting was found for the CDW transition in a one-dimensional model for spinless fermions with nearest-neighbor interactions [24]. Apparently, much larger systems are required to overcome this systematic error. As in Ref. [23], we use the critical interaction strengths $U_{\rm c,LL}(V)$ as an upper bound to the exact value $U_{\rm c}(V)$.
IV.4 Transitions at fixed Hubbard interaction
However, achieving the theoretical value of $K=1$ within the metallic Luttinger liquid is hindered by finite-size effects. Note that the results from Fig. 13 leads to the vertical lines for $U=1.3$ in the phase diagram.
@@ IV.4 Transitions at fixed Hubbard interaction
Of course, it is not possible to specify the nature of the insulating phases from the Luttinger parameter alone. One might argue that, due to finite-size effects, the phase transition into the CDW phase is more easily detected, i.e., the curve $K(U,V)$ for fixed $U$ is steeper as a function of $V$ in the CDW phase than in the Mott insulating phase. This ambiguity does not pose a problem because other physical quantities, e.g., the structure factor, permit to identify the CDW phase, as we discuss next.
@@ IV.6.2 Fano resonance curve
All approaches lead to consistent results. The extrapolations from the Luttinger exponent underestimate the critical interactions whereas the results from the Fano resonance give an upper bound to $V_{\rm c}^{\rm CDW}(U)$. As for the Mott transition, this permits to set boundaries to the exact critical value of the metal-to-insulator transition.
@@ V.1 Quantum phase diagrams
For convenience, we reproduce the quantum phase diagrams in Fig. 1 in Fig. 17. As in our previous work [23], we refrain from the analysis of the transition between Mott and CDW insulators. As shown numerically for the standard extended Hubbard model [30, 31, 32, 33], the transition is discontinuous for large interactions, $U,V\gg W$, as also inferred from the atomic limit, see appendix. For moderate interactions, $U,V\approx W$, and fixed $U$, a transition occurs from the Mott insulator to a bond-order-wave insulator as a function of $V$ before the CDW insulator dominates at large nearest-neighbor interactions.
Beyond the triple point where Luttinger liquid, Mott insulator, and CDW insulator meet in our phase diagrams, we also expect to find a region with a bond-order wave in extended $1/r$-Hubbard models. Therefore, the dashed line separating Mott-Hubbard insulator and charge-density-wave insulator should be taken as a guide to the eye only. For this reason, we only show the phase diagram for $U\leq 1.9$, and the analysis of the triple point and the bond-order wave phase are beyond the scope of our present work.
@@
Figure 17: (a) Phase diagram of the one-dimensional $1/r$-Hubbard model with nearest-neighbor interactions. (b) Phase diagram of the one-dimensional $1/r$-Hubbard model with $1/r$-long-range interactions. Continuous lines: second order polynomial fits; dotted line: Hartree-Fock (HF) result for the transition between metal and charge-density-wave insulator.
V.1.1 Parametrization of the CDW critical line
@@
V.1.2 Parametrization of the Mott critical line
@@
where $b_{i}$ naturally differ for nearest-neighbor and $1/r$-long-range interactions. The corresponding critical lines, $V_{\rm c}^{\rm Mott}(U)$, are shown in the phase diagrams 17. In the range $1\leq U\leq 1.9$ and due to the fact that $b_{3}$ is small, we can approximate
@@ VI.2.2 Advantages for numerical investigations
VI.2.3 Remaining issues

### v2

@@ I.2 Phase diagrams
Figure 1: (a) Phase diagram of the one-dimensional $1/r$-Hubbard model with nearest-neighbor interactions $V_{\rm NN}$. (b) Phase diagram of the one-dimensional $1/r$-Hubbard model with $1/r$-long-range interactions. Dots: estimate for the critical interaction, $\overline{U}_{\rm c}$, with error bounds; continuous lines: polynomial fits, see Sect. V.
For dominant Coulomb interactions, the separation line between the Mott-Hubbard insulator and the CDW insulator is given by $V_{\rm NN}=U/2$ to leading order, with a stabilization tendency toward the Mott-Hubbard insulator in the next order in $1/U$.
In the absence of inter-site interactions, the Mott-Hubbard transition occurs at $U_{\rm c}(V=0)=W$ [20, 6, 22]. The inclusion of a repulsive nearest-neighbor interaction stabilizes the metallic phase because the additional repulsive interaction softens the two-particle scattering potential in position space.
@@
On the other hand, in the atomic limit, the CDW and Mott-Hubbard insulators are separated by $V_{\rm c}^{\rm CDW}(U)=U/(2\ln(2))\approx 0.721U$ for the $1/r$-Hubbard model with $1/r$-long-range interactions, see appendix. Thus, the renormalization factor in the atomic limit is given by $R^{\rm al}_{\rm Mott}=(1/2)/[1/(2\ln 2)]=\ln 2\approx 0.693$. In Fig. 2, the Mott critical lines $U_{\rm c}^{\rm Mott}(V)$ of the $1/r$-Hubbard models with nearest-neighbor and $1/r$-long-range interactions are then mapped onto each other using $V_{\rm NN}=R_{\rm Mott}V_{\rm LR}$, $R_{\rm Mott}=0.69$.
@@
Before we proceed, we caution the reader that our findings relate to the ground-state and low-energy physics of interacting quantum particles. It is well known that an attractive potential well can be used to describe semi-quantitatively the $1s$ bound state of the hydrogen atom or of an exciton. However, such an approach fails to describe properly the fine structure of the wave functions or the energy of higher-excited states. Similarly, the long-range tail of the Coulomb interaction can give rise to exotic phenomena, particularly under non-equilibrium conditions. Long-range interactions are also more significant in dilute bosonic systems, see the example of polar bosonic gases in one-dimensional lattice [24], or in the case of bosonic many-body localization [25], and references therein.
@@ II.2.3 Density-density correlation function and Luttinger parameter
We extract the Luttinger parameter $K(U,V)$ from the structure factor,
@@
where we denote the structure factor in the thermodynamic limit by $\tilde{C}^{\rm NN}(q;U,V)$. Using Eq. (30), the Luttinger parameter can be calculated numerically with very good accuracy [31]. The Luttinger parameter can be used to locate the metal-insulator transition in one dimension.
@@ III.1 Ground-state energy
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
Figure 4: Ground-state energy density at half band-filling, $e_{0}(n=1;U,V)$, for the $1/r$-Hubbard model with $1/r$-long-range interactions (crosses) from the extrapolation to the thermodynamic limit in Fig. 3, and for the extended $1/r$-Hubbard model (open triangles, see Ref. [23]). The continuous line is the exact result for $V=0$, $e_{0}(n=1;U,V=0)$.
@@ III.2 Two-particle gap
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
In the bare $1/r$-Hubbard model, the Mott transition is solely governed by the on-site interaction. Once the Hubbard interaction exceeds the bandwidth, $U>W$, electrons become ‘localized’, and the metallic state goes over to a Mott insulator. In the presence of inter-site interactions, the Mott transition is shifted to larger values which implies a more stable metallic Luttinger liquid.
At first glance, the shift of the Mott transition to higher values in models with repulsive inter-site interactions seems counterintuitive. From a wave-mechanical viewpoint, however, the repulsive interaction between lattice sites softens the two-particle scattering potential. Figuratively speaking, particles scattered by the weaker interaction $V$ do not fully experience the stronger on-site interaction $U$. This ties in with the observation that the gap in the $1/r$-Hubbard model with $1/r$-long-range interactions is smaller than that for the extended $1/r$-Hubbard model for the same $v$ because the long-range parts of the inter-site interaction further smoothen the two-particle scattering potential in position space.
@@
III.3 Momentum distribution
These considerations show that the Luttinger parameter must deviate from unity, $K(U,V)<1$, for all $(U,V)$, even though corrections to unity are (exponentially) small for $U,V\ll 1$. Therefore, $n(k)$ is a continuous function in the $1/r$-Hubbard model for all $U,V>0$.
III.4 Structure factor and CDW order parameter
@@ IV.3 Luttinger parameter
As seen from Table 1, the critical value from the Luttinger parameter systematically overestimates the correct interaction strengths by some five percent. The same overshooting was found for the CDW transition in a one-dimensional model for spinless fermions with nearest-neighbor interactions [26].
Apparently, much larger systems are required to overcome this systematic error. As in Ref. [23], we use the critical interaction strengths $U_{\rm c,LL}(V)$ as an upper bound to the exact value $U_{\rm c}(V)$.
@@
IV.4 Transitions at fixed Hubbard interaction
However, achieving the theoretical value of $K=1$ within the metallic Luttinger liquid is hindered by finite-size effects. Note that the results from Fig. 13 lead to the vertical lines for $U=1.3$ in the phase diagram.
Of course, it is not possible to specify the nature of the insulating phases from the Luttinger parameter alone. One might argue that, due to finite-size effects, the phase transition into the CDW phase is more easily detected, i.e., the curve $K(U,V)$ for fixed $U$ is steeper as a function of $V$ in the CDW phase than in the Mott insulating phase.
This ambiguity does not pose a problem because other physical quantities, e.g., the structure factor, permit to identify the CDW phase, as we discuss next.
@@ IV.6.2 Fano resonance curve
All approaches lead to consistent results. The extrapolations from the Luttinger parameter underestimate the critical interactions whereas the results from the Fano resonance give an upper bound to $V_{\rm c}^{\rm CDW}(U)$. As for the Mott transition, this permits to set boundaries to the exact critical value of the metal-to-insulator transition.
@@ V.1 Quantum phase diagrams
For convenience, we reproduce the quantum phase diagrams in Fig. 1 in Fig. 17, where we add the Hartree-Fock results for the transition between metal and CDW insulator as dotted lines and the strong-coupling limit for the separation lines between Mott-Hubbard and CDW insulators as dashed lines, respectively. As in our previous work [23], we refrain from the analysis of the transition between Mott and CDW insulators.
@@
Figure 17: (a) Phase diagram of the one-dimensional $1/r$-Hubbard model with nearest-neighbor interactions. (b) Phase diagram of the one-dimensional $1/r$-Hubbard model with $1/r$-long-range interactions. Continuous lines: polynomial fits of the phase-separation lines, see Sects. V.1.1 and V.1.2; dotted line: Hartree-Fock (HF) result for the transition between metal and charge-density-wave insulator.
As shown numerically for the standard extended Hubbard model [32, 33, 34, 35], the transition is discontinuous for large interactions, $U,V\gg W$, as also inferred from the atomic limit, see appendix, where $V_{\rm NN,c}^{\rm al}=U/2$ for the extended Hubbard model. Since the kinetic energy slightly stabilizes the Mott-Hubbard insulator, the phase separation line in Fig. 17a for large interactions is approximately given by $V_{\rm c,NN}(U\gg W)\approx 0.51U$ (dashed line). The same line of reasoning gives $V_{\rm c,LR}^{\rm al}=U/[2\ln(2)]$ in the atomic limit in the presence of $1/r$-long-range interactions, and $V_{\rm c,LR}(U\gg W)\approx 0.733U\approx 0.51U/\ln(2)$ (dashed line).
For the standard extended Hubbard model and for moderate interactions, $U,V\approx W$, and fixed $U$, a transition occurs from the Mott insulator to a bond-order-wave insulator as a function of $V$ before the CDW insulator dominates at large nearest-neighbor interactions. Beyond the triple point where Luttinger liquid, Mott insulator, and CDW insulator meet in our phase diagrams, we also expect to find a region with a bond-order wave in extended $1/r$-Hubbard models. Therefore, the dashed line separating Mott-Hubbard insulator and charge-density-wave insulator should be taken as a guide to the eye only. For this reason, we only show the phase diagram for $U\leq 1.9$, and the analysis of the triple point and the bond-order wave phase are beyond the scope of our present work.
V.1.1 Parameterization of the CDW critical line
@@
V.1.2 Parameterization of the Mott critical line
@@
where $b_{i}$ naturally differ for nearest-neighbor and $1/r$-long-range interactions. The corresponding critical lines, $V_{\rm c}^{\rm Mott}(U)$, are shown in the phase diagrams 17. In the range $1\leq U\leq 1.9$ and due to the fact that $b_{3}$ is small enough, we can approximate
@@ VI.2.2 Advantages for numerical investigations
VI.2.3 Correlation functions
Models with algebraically decaying interactions may display algebraically decaying correlation functions even in the presence of a gap, see, e.g., Ref. [43]. Therefore, it would be interesting to study in detail the behavior of real-space correlation functions close to the quantum phase transitions.
Unfortunately, as seen from Figs. 7 and 8, our systems sizes $L\leq 64$ are too small to address the asymptotic region $L\gg r\gg 1$ where short-range correlations have died out $(r\gg 1$) and finite-size effects are still negligible, $r\ll L$. The Fourier transform of the single-particle density matrix, the momentum distribution $n(k)$, looks discontinuous at first glance despite the fact that it is continuous both in the Luttinger liquid and in the Mott and CDW insulating phases. Likewise, the density-density correlation function $\tilde{C}^{\rm NN}(q)$ looks unspectacular and it is difficult, yet feasible, to extract the Luttinger parameter from it using eq. (30). Much larger systems, of the order of ten thousand sites, are required for a detailed analysis of the asymptotic behavior of real-space correlation functions. Fortunately, for our analysis of the quantum phase diagram, we do not have to rely on correlation functions.
VI.2.4 Remaining issues